# Flow control of three-dimensional cylinders transitioning to turbulence via multi-agent reinforcement learning

### v1

@@ Introduction
The transportation industry, and the aerospace sector in particular, require new ground-breaking methods to overcome the challenges that they are currently facing, i.e. the need to reduce fossil-fuel-related emissions. The implementation of flow-control systems, both passive and active, plays a vital role in the development of more sustainable solutions that can drastically reduce fuel usage, mitigate air and noise pollution, and even improve maneuverability [1]. The aerodynamic drag in subsonic aircraft is divided mainly into pressure, skin friction (due to viscous stresses), and lift-induced components. Wing-tip effects aside, the two former are the dominant terms.
@@
Control devices utilize aerodynamic principles to manipulate pressure and viscosity, effectively reducing drag. For instance, slats and flaps are control surfaces located on the leading and trailing edges of an airfoil which impact the aircraft operational conditions [2]. Modern advancements include winglets [3] aimed at mitigating the lift-induced drag or vortex generators [4], which are used to control boundary-layer separation. Such developments have significantly improved aerodynamic performance. Additionally, a number of alternative approaches like morphing surfaces, spiroids, or blowing devices [5, 6] are currently being explored. Despite their extensive potential, designing optimal geometries or strategies for these devices has raised significant challenges due to the massive computational resources required to tackle such an intricate interplay between pressure and viscous effects in all flight regimes. Nevertheless, ongoing research efforts and current computational innovations are leading to further advancements toward achieving optimal global control.
@@
In parallel with the recent innovations in flow control, the irruption of machine-learning (ML) techniques has brought tremendous potential to the aeronautics industry, both in terms of studying fundamental problems in fluid mechanics [7, 8] and devising completely new strategies for active and passive flow control (AFC and PFC, respectively) [9]. Deep reinforcement learning (DRL) is one of the fastest-growing fields within ML  [10] and one of the techniques attracting most interest. Expanding on its success in tabletop games [11], DRL works well in any system where a controller interacts with an environment to improve a task. That is the case for most AFC cases since DRL interacts with the flow on the fly and receives feedback from it, gaining experience and progressively improving the choice of actions.
@@
AFC setups are complex high-dimensional problems that require substantial computational resources to find the optimal values within the large parametric space of the control system. DRL and neural networks have emerged as valuable tools to make this process feasible, enabling the development of effective control strategies at a reasonable computational cost. The literature on DRL for AFC grows at a fast pace, exhibiting studies on flow control for two-dimensional (2D) cylinders ranging from $Re_{D}=100$ and $2000$ (where $Re_{D}$ is the Reynolds number based on inflow velocity $U_{\infty}$ and cylinder diameter $D$) with $17\%$ and $38\%$ drag reduction, respectively [12, 13, 14, 15, 16, 17, 18], aircraft wings [19], fluid-structure interaction [20], turbulent channels [21], shape optimization [22, 23, 24], Rayleigh–Bénard convection [25] or turbulence modelling [26, 27, 28, 29]. Some recent literature demonstrates the possibility of transfer learning from exploration done in 2D cylinders to 3D domains and higher $Re_{D}$: in Ref. [30] the wake of a cylinder is controlled by means of two rotating cylinders and in Ref. [18] the control is carried out through multiple jets over the cylinder surface. The present work extends this state-of-the-art in 3D cylinders, considering multiple actuators governed by the novel implementation of a multi-agent reinforcement learning (MARL) framework into a setup based on a distributed-input distributed-output (DIDO) scheme. In our case, the agent focuses on exploring the underlying 3D physics during training, and the AFC is implemented by multiple independent zero-net-mass-flow (ZNMF) jets placed along the cylinder span and aligned along two slots on the top and bottom surfaces. To the best of the authors’ knowledge, this work marks the first time where exploration sessions are directly conducted within 3D cylinders.
@@
As the Reynolds number increases, the flow around a cylinder exhibits different characteristics. Initially, up to approximately $Re_{D}\approx 40$, steady laminar flow prevails, characterized by symmetric counter-rotating vortices in the near wake. Beyond $Re_{D}\approx 190$, laminar vortex shedding emerges, forming the well-known Kármán vortex street. In the subsequent regimes, between $260190<Re_{D}<$ 260, the mode-A instability, characterized by dominant spanwise wavelengths of $\lambda_{z}=4D$ [31, 32] is dominant. As $Re_{D}\approx 260$ is surpassed, mode B becomes predominant, and finer three-dimensional features with shorter wavelengths of $\lambda_{z}=1D$ are found. Beyond these regimes, the cylinder wake evolves into a more chaotic and turbulent state.
@@
Discovering flow-control strategies for the flow around a cylinder when the wake transitions from 2D to 3D is challenging. The MARL setup needs to exploit the characteristics of the spanwise structures as the wake becomes three-dimensional to devise effective control approaches. The transition range of for 2D environments has been shown to be suitable, showcasing the generalization ability of deep neural networks [12]. However, it has been widely recognized that studying $Re_{D}>250$ in a 2D context leads to inaccurate predictions of aerodynamic forces. In the present work, the exploration of the 3D context allows tackling possible novel strategies that take advantage of the drag reduction originated by 3D instabilities.
@@
DRL is based on maximizing a reward ($R$) provided to an agent interacting continuously with an environment through actions $A$. The agent receives information about the environment state at each actuation step through partial observations $S$ of the system. A sequence of consecutive actions is denoted as an episode. When a batch of episodes is finished, the agent updates the neural-network weights to progressively determine a policy that maximizes the expected reward for a given $S$. For a detailed understanding of the most recent advances in flow control with MARL we refer to Refs. [33, 34].
@@ Results
We also study the ratio between the total fluid mass intercepted by the frontal area of the cylinder $E_{\infty}$ and the total mass used by the actuators $E_{c}$. Based on the definitions used in Ref. [18], we propose the following expression for the ratio $E_{c}^{*}$:
@@
Figure 4: Visualizations of the flow coherent structures during DRL-based strategy exploitation. Temporal evolution of the flow at (a) $Re_{D}=300$ and (b) $Re_{D}=400$, from uncontrolled state to a stable DRL control. (Top) Non-dimensional mass-flow rate per unit width $Q$, lift coefficient $C_{l}$ and drag coefficient $C_{d}$ as a function of time; (bottom) snapshots showing vortical structures identified with the $\lambda_{2}$ criterion [35], where the isosurface $\lambda_{2}D^{2}/U_{\infty}^{2}=-0.5$ is shown for uncontrolled, transient, and DRL-controlled states. The colors framing the flow visualizations correspond to the instants indicated in the temporal evolution of the relevant flow quantities.
@@ Methods
where $Q=\dot{m}/L_{z}$ and $|\theta-\theta_{0}|\in[-\omega/2,\omega/2]$, $\dot{m}$ is the mass flow rate. The absolute value of the jet velocity is projected into the $x$ and $y$ axes, since $\lVert{U_{\rm{jet}}}\lVert$ corresponds to the radial cylinder direction. For each pseudo-environment, we set opposite action values within the pair of top and bottom jets i.e. $Q_{\rm{90^{\circ}}}=-Q_{\rm{270^{\circ}}}$, in order to ensure the global zero net mass flux. An earlier version of this setup was developed in Ref. [36].
A conceptually similar control approach was reported in Ref. [37], in particular the one they denote as out-phase. Out-phase approach consists of different constant mass-flow sinusoidal distributions along the spanwise with different wavelengths.
@@
The transition from laminar to the emergence of first three-dimensional instabilities in the cylinder wake occurs within a range of $Re_{D}=200$ and $300$ [31, 38, 39, 40, 41, 42]. The motivation of this research is to challenge the control system to discover the optimal strategies as the flow starts to become three-dimensional. By being able to manipulate scales with different spanwise wavelengths in the wake, the DRL-based control can learn effective mechanisms leading to a substantial reduction of the drag in the cylinder.
The numerical simulations are carried out by means of the numerical solver Alya, which is described in detail in Ref. [43]. The spatial discretization is based on the finite-element method (FEM) and the incompressible Navier–Stokes equations:
@@
are integrated numerically, where $\bm{\epsilon}$ is a function of the velocity $\bm{u}$ which defines the velocity strain-rate tensor $\bm{\epsilon}=1/2(\nabla\bm{u}+(\nabla\bm{u})^{\mathrm{T}})$ and $\bm{f}$ are the external body forces. In Equation (4), the convective term $(\bm{u}\cdot\nabla)\bm{u}$ is expressed as a term conserving energy, momentum and angular momentum [44, 45]. For the time discretization, a semi-implicit method is used where the convective term follows a second-order Runge–Kutta scheme and a Crank–Nicholson scheme is used for the diffusive term [46]. To select the appropriate time step, Alya uses an eigenvalue-based time-integration scheme [47]. Then, for each time step, the numerical solution of these equations is computed. Drag and lift forces ($F_{x}$ and $F_{y}$, respectively) forces are obtained by integration over the cylinder surface $\bm{s}$:
@@
In the present work, we implemented a deep-reinforcement-learning (DRL) framework using Tensorforce libraries [48]. DRL is very well suited for unsteady flow-control problems as it provides the possibility to dynamically interact with an environment, being able to dynamically set the actuation based on the varying flow state. We use the proximal-policy-optimization (PPO) algorithm [49], which is a policy-gradient approach based on a surrogate loss function for policy updates to prevent drastic drops in performance. This algorithm demonstrates robustness, as it is somewhat forgiving with hyperparameter initializations and can perform adequately across a diverse range of RL tasks without extensive tuning.
The neural-network architecture consists of two dense layers of 512 neurons each. The batch size, i.e. the total number of experiences that the PPO agent uses for each gradient-descent iteration, is set to 80, which is larger than the typical values used in 2D trainings [50, 51], but sufficiently small to efficiently update the neural-network parameters. If there is a discrepancy between the batch size and the environments running at the same time there is a risk of wasting information that will not be captured by the agent. The limitation is that we have 10 actuators per environment and we need 10 streamed experiences which will be synchronized, so we have to work with a total of $10n_{\rm environments}$ set of experiences. A streamed experience consists of a set of states, actions, rewards, and the predicted state that the agent expects to achieve. It is denoted as $(S,A,R,S^{\prime})_{i,t}$ for each pseudo-environment, and each of the Reynolds numbers under consideration has its own agent and policy.
@@
The interactions between the agent and the physical environment are denoted as actions $A$, and they influence the system during $T_{a}$ time units. We update the jet boundary conditions using Equation (2). The shift in time between actions, $Q_{t}\rightarrow Q_{t+1}$ is managed through an exponential function. The smooth transition diminishes the appearance of sudden discontinuities which can spoil a training process. The DRL library outputs values in the range of $[-1,1]$, requiring rescaling as $Q=AQ_{\rm{max}}$ to avoid excessively large actuations. Hence, $Q_{\rm{max}}=0.176$ was set based on our experience with DRL for flow control, and corresponds to twice the values used in the 2D cylinder setups [50].
Certain parameters in the DRL configuration are closely tied to the fluid-mechanics problem under consideration. The episode duration is specifically defined to include at least six vortex-shedding periods ($T_{k}=1/f_{k}$). We set $T_{a}<0.05T_{k}$, based on the experience gathered with previous studies [12, 51]. This allows sufficient time between actions to produce an effect on the flow. Note that if the time between actions is too short, there will be noise in the training process and it will become difficult to converge. On the other hand, if this is too large the agent will not be able to control the smaller-scale structures associated with shorter time scales. Thus, a total of 120 actuations per episode is deemed sufficient for evaluating the cumulative reward. It is noteworthy that each episode starts from an uncontrolled converged state of the problem. This corresponds to what happens during training, but when we evaluate the DRL model in exploitation mode (also denoted as a deterministic mode), we make the episodes 4 times longer to ensure statistical convergence.
@@
Extensive work documented in Ref. [36] was carried out to adjust the MARL framework and the communications setup. For instance, the definition of $S$ is the result of a compromise between computational practicality and physical relevance. The spanwise wavelength of the structures in the wake also helped to define the spacing of the $xy$ planes defining the system state. Note that the number of data points used for this state correlates with the number of weights calculated for the first fully connected layer of the neural network.

### v2

@@ Introduction
The transportation industry, and the aerospace sector in particular, require new ground-breaking methods to overcome the challenges that they are currently facing, i.e. the need to reduce fossil-fuel-related emissions. The implementation of flow-control systems, both passive and active, plays a vital role in the development of more sustainable solutions that can drastically reduce fuel usage, mitigate air and noise pollution, and even improve maneuverability [choi_review_2008]. The aerodynamic drag in subsonic aircraft is divided mainly into pressure, skin friction (due to viscous stresses), and lift-induced components. Wing-tip effects aside, the two former are the dominant terms.
@@
Control devices utilize aerodynamic principles to manipulate pressure and viscosity, effectively reducing drag. For instance, slats and flaps are control surfaces located on the leading and trailing edges of an airfoil which impact the aircraft operational conditions [raymer_book_2012]. Modern advancements include winglets [whitcomb_1976] aimed at mitigating the lift-induced drag or vortex generators [lin_vortex_2002], which are used to control boundary-layer separation. Such developments have significantly improved aerodynamic performance. Additionally, a number of alternative approaches like morphing surfaces, spiroids, or blowing devices [siddiqui_2017, guerrero_2012] are currently being explored. Despite their extensive potential, designing optimal geometries or strategies for these devices has raised significant challenges due to the massive computational resources required to tackle such an intricate interplay between pressure and viscous effects in all flight regimes. Nevertheless, ongoing research efforts and current computational innovations are leading to further advancements toward achieving optimal global control.
@@
In parallel with the recent innovations in flow control, the irruption of machine-learning (ML) techniques has brought tremendous potential to the aeronautics industry, both in terms of studying fundamental problems in fluid mechanics [vinuesa_transformative_2023, vinuesa_enhancing_2022] and devising completely new strategies for active and passive flow control (AFC and PFC, respectively) [Le_Clainche_aircraft_2023]. Deep reinforcement learning (DRL) is one of the fastest-growing fields within ML  [garnier_review_2021] and one of the techniques attracting most interest. Expanding on its success in tabletop games [alphago_2016], DRL works well in any system where a controller interacts with an environment to improve a task. That is the case for most AFC cases since DRL interacts with the flow on the fly and receives feedback from it, gaining experience and progressively improving the choice of actions.
@@
AFC setups are complex high-dimensional problems that require substantial computational resources to find the optimal values within the large parametric space of the control system. DRL and neural networks have emerged as valuable tools to make this process feasible, enabling the development of effective control strategies at a reasonable computational cost. The literature on DRL for AFC grows at a fast pace, exhibiting studies on flow control for two-dimensional (2D) cylinders ranging from $Re_{D}=100$ and $2000$ (where $Re_{D}$ is the Reynolds number based on inflow velocity $U_{\infty}$ and cylinder diameter $D$) with $17\%$ and $38\%$ drag reduction, respectively [tang_robust_2020, xu_active_2020, paris_robust_2021, LiZhang_2022, ren_applying_2021, dixia_fan_bluffexperiments_2020, chatzimanolakis2023drag], aircraft wings [vinuesa_flow_2022], fluid-structure interaction [chen_deep_2023], turbulent channels [guastoni_deep_2023], shape optimization [yan_2019, viquerat2021direct, keramati_2022], Rayleigh–Bénard convection [vignon_effective_2023] or turbulence modelling [kurz_modelling_2023, novati_2021, beck_toward_2023, bae2022scientific]. Some recent literature demonstrates the possibility of transfer learning from exploration done in 2D cylinders to 3D domains and higher $Re_{D}$: in Ref. [Wang_Karniadakis_2023] the wake of a cylinder is controlled by means of two rotating cylinders and in Ref. [chatzimanolakis2023drag] the control is carried out through multiple jets over the cylinder surface. The present work extends this state-of-the-art in 3D cylinders, considering multiple actuators governed by the novel implementation of a multi-agent reinforcement learning (MARL) framework into a setup based on a distributed-input distributed-output (DIDO) scheme. In our case, the agent focuses on exploring the underlying 3D physics during training, and the AFC is implemented by multiple independent zero-net-mass-flow (ZNMF) jets placed along the cylinder span and aligned along two slots on the top and bottom surfaces. To the best of the authors’ knowledge, this work marks the first time where exploration sessions are directly conducted within 3D cylinders.
@@
As the Reynolds number increases, the flow around a cylinder exhibits different characteristics. Initially, up to approximately $Re_{D}\approx 40$, steady laminar flow prevails, characterized by symmetric counter-rotating vortices in the near wake. Beyond $Re_{D}\approx 190$, laminar vortex shedding emerges, forming the well-known Kármán vortex street. In the subsequent regimes, between $260190<Re_{D}<$ 260, the mode-A instability, characterized by dominant spanwise wavelengths of $\lambda_{z}=4D$ [williamson_vortex_1996, Barkley_Henderson_1996] is dominant. As $Re_{D}\approx 260$ is surpassed, mode B becomes predominant, and finer three-dimensional features with shorter wavelengths of $\lambda_{z}=1D$ are found. Beyond these regimes, the cylinder wake evolves into a more chaotic and turbulent state.
@@
Discovering flow-control strategies for the flow around a cylinder when the wake transitions from 2D to 3D is challenging. The MARL setup needs to exploit the characteristics of the spanwise structures as the wake becomes three-dimensional to devise effective control approaches. The transition range of for 2D environments has been shown to be suitable, showcasing the generalization ability of deep neural networks [tang_robust_2020]. However, it has been widely recognized that studying $Re_{D}>250$ in a 2D context leads to inaccurate predictions of aerodynamic forces. In the present work, the exploration of the 3D context allows tackling possible novel strategies that take advantage of the drag reduction originated by 3D instabilities.
@@
DRL is based on maximizing a reward ($R$) provided to an agent interacting continuously with an environment through actions $A$. The agent receives information about the environment state at each actuation step through partial observations $S$ of the system. A sequence of consecutive actions is denoted as an episode. When a batch of episodes is finished, the agent updates the neural-network weights to progressively determine a policy that maximizes the expected reward for a given $S$. For a detailed understanding of the most recent advances in flow control with MARL we refer to Refs. [brunton_closed-loop_2015, vignon_recent_2023].
@@ Results
We also study the ratio between the total fluid mass intercepted by the frontal area of the cylinder $E_{\infty}$ and the total mass used by the actuators $E_{c}$. Based on the definitions used in Ref. [chatzimanolakis2023drag], we propose the following expression for the ratio $E_{c}^{*}$:
@@
Figure 4: Visualizations of the flow coherent structures during DRL-based strategy exploitation. Temporal evolution of the flow at (a) $Re_{D}=300$ and (b) $Re_{D}=400$, from uncontrolled state to a stable DRL control. (Top) Non-dimensional mass-flow rate per unit width $Q$, lift coefficient $C_{l}$ and drag coefficient $C_{d}$ as a function of time; (bottom) snapshots showing vortical structures identified with the $\lambda_{2}$ criterion [Jeong_Hussain_1995], where the isosurface $\lambda_{2}D^{2}/U_{\infty}^{2}=-0.5$ is shown for uncontrolled, transient, and DRL-controlled states. The colors framing the flow visualizations correspond to the instants indicated in the temporal evolution of the relevant flow quantities.
@@ Methods
where $Q=\dot{m}/L_{z}$ and $|\theta-\theta_{0}|\in[-\omega/2,\omega/2]$, $\dot{m}$ is the mass flow rate. The absolute value of the jet velocity is projected into the $x$ and $y$ axes, since $\lVert{U_{\rm{jet}}}\lVert$ corresponds to the radial cylinder direction. For each pseudo-environment, we set opposite action values within the pair of top and bottom jets i.e. $Q_{\rm{90^{\circ}}}=-Q_{\rm{270^{\circ}}}$, in order to ensure the global zero net mass flux. An earlier version of this setup was developed in Ref. [suarez2023active].
A conceptually similar control approach was reported in Ref. [choi_distributed_2005], in particular the one they denote as out-phase. Out-phase approach consists of different constant mass-flow sinusoidal distributions along the spanwise with different wavelengths.
@@
The transition from laminar to the emergence of first three-dimensional instabilities in the cylinder wake occurs within a range of $Re_{D}=200$ and $300$ [williamson_vortex_1996, bays1993streamwise, bloor1964transition, karniadakis1989frequency, karniadakis1992three, norberg1994experimental]. The motivation of this research is to challenge the control system to discover the optimal strategies as the flow starts to become three-dimensional. By being able to manipulate scales with different spanwise wavelengths in the wake, the DRL-based control can learn effective mechanisms leading to a substantial reduction of the drag in the cylinder.
The numerical simulations are carried out by means of the numerical solver Alya, which is described in detail in Ref. [vazquez_alya_2016]. The spatial discretization is based on the finite-element method (FEM) and the incompressible Navier–Stokes equations:
@@
are integrated numerically, where $\bm{\epsilon}$ is a function of the velocity $\bm{u}$ which defines the velocity strain-rate tensor $\bm{\epsilon}=1/2(\nabla\bm{u}+(\nabla\bm{u})^{\mathrm{T}})$ and $\bm{f}$ are the external body forces. In Equation (4), the convective term $(\bm{u}\cdot\nabla)\bm{u}$ is expressed as a term conserving energy, momentum and angular momentum [charnyi_conservation_2017, charnyi_efficient_2019]. For the time discretization, a semi-implicit method is used where the convective term follows a second-order Runge–Kutta scheme and a Crank–Nicholson scheme is used for the diffusive term [crank_practical_nodate]. To select the appropriate time step, Alya uses an eigenvalue-based time-integration scheme [article_trias]. Then, for each time step, the numerical solution of these equations is computed. Drag and lift forces ($F_{x}$ and $F_{y}$, respectively) forces are obtained by integration over the cylinder surface $\bm{s}$:
@@
In the present work, we implemented a deep-reinforcement-learning (DRL) framework using Tensorforce libraries [schaarschmidt_tensorforce_2017]. DRL is very well suited for unsteady flow-control problems as it provides the possibility to dynamically interact with an environment, being able to dynamically set the actuation based on the varying flow state. We use the proximal-policy-optimization (PPO) algorithm [schulman2017proximal], which is a policy-gradient approach based on a surrogate loss function for policy updates to prevent drastic drops in performance. This algorithm demonstrates robustness, as it is somewhat forgiving with hyperparameter initializations and can perform adequately across a diverse range of RL tasks without extensive tuning.
The neural-network architecture consists of two dense layers of 512 neurons each. The batch size, i.e. the total number of experiences that the PPO agent uses for each gradient-descent iteration, is set to 80, which is larger than the typical values used in 2D trainings [varela_deep_2022, rabault_artificial_2019], but sufficiently small to efficiently update the neural-network parameters. If there is a discrepancy between the batch size and the environments running at the same time there is a risk of wasting information that will not be captured by the agent. The limitation is that we have 10 actuators per environment and we need 10 streamed experiences which will be synchronized, so we have to work with a total of $10n_{\rm environments}$ set of experiences. A streamed experience consists of a set of states, actions, rewards, and the predicted state that the agent expects to achieve. It is denoted as $(S,A,R,S^{\prime})_{i,t}$ for each pseudo-environment, and each of the Reynolds numbers under consideration has its own agent and policy.
@@
The interactions between the agent and the physical environment are denoted as actions $A$, and they influence the system during $T_{a}$ time units. We update the jet boundary conditions using Equation (2). The shift in time between actions, $Q_{t}\rightarrow Q_{t+1}$ is managed through an exponential function. The smooth transition diminishes the appearance of sudden discontinuities which can spoil a training process. The DRL library outputs values in the range of $[-1,1]$, requiring rescaling as $Q=AQ_{\rm{max}}$ to avoid excessively large actuations. Hence, $Q_{\rm{max}}=0.176$ was set based on our experience with DRL for flow control, and corresponds to twice the values used in the 2D cylinder setups [varela_deep_2022].
Certain parameters in the DRL configuration are closely tied to the fluid-mechanics problem under consideration. The episode duration is specifically defined to include at least six vortex-shedding periods ($T_{k}=1/f_{k}$). We set $T_{a}<0.05T_{k}$, based on the experience gathered with previous studies [tang_robust_2020, rabault_artificial_2019]. This allows sufficient time between actions to produce an effect on the flow. Note that if the time between actions is too short, there will be noise in the training process and it will become difficult to converge. On the other hand, if this is too large the agent will not be able to control the smaller-scale structures associated with shorter time scales. Thus, a total of 120 actuations per episode is deemed sufficient for evaluating the cumulative reward. It is noteworthy that each episode starts from an uncontrolled converged state of the problem. This corresponds to what happens during training, but when we evaluate the DRL model in exploitation mode (also denoted as a deterministic mode), we make the episodes 4 times longer to ensure statistical convergence.
@@
Extensive work documented in Ref. [suarez2023active] was carried out to adjust the MARL framework and the communications setup. For instance, the definition of $S$ is the result of a compromise between computational practicality and physical relevance. The spanwise wavelength of the structures in the wake also helped to define the spacing of the $xy$ planes defining the system state. Note that the number of data points used for this state correlates with the number of weights calculated for the first fully connected layer of the neural network.